OpenStudy (el_arrow):

limit (x^2-x^3)e^2x x-->-infinity

OpenStudy (el_arrow):

this is what i got so far

OpenStudy (el_arrow):

(x^2-x^3)/e^-2x

OpenStudy (el_arrow):

i later did the derivative which is (2x-3x^2)/-2e^-2x

OpenStudy (el_arrow):

is it right?

OpenStudy (el_arrow):

my book says that the answer is 0

zepdrix (zepdrix):

So you were able to get an indeterminate form of \(\Large\rm \frac{\infty}{\infty}\) which allowed you to apply L'Hospital's Rule? Yes, looks good so far :) Derivative is correct.

OpenStudy (el_arrow):

i am having trouble find the answer

zepdrix (zepdrix):

What form are you getting now? Is it still indeterminate?

OpenStudy (el_arrow):

yes so now its 2-6x/4e^-2x

OpenStudy (el_arrow):

right?

zepdrix (zepdrix):

Good good good. Notice that the x's are deteriorating while the exponential is super strong. Can't touch that guy :) So you will EVENTUALLY get to a point where only the exponential is left. You might have to L'Hop one more time.

OpenStudy (el_arrow):

so then i got -6/-8e^-2x

zepdrix (zepdrix):

Just make sure you're checking the form between each application of L'Hop. It has to stay in the form \(\Large\rm \frac{0}{0}\) or \(\Large\rm \frac{\infty}{\infty}\)

zepdrix (zepdrix):

mmm k

OpenStudy (el_arrow):

so when i plug in the -infinity times the -2 do i get positive infinity?

zepdrix (zepdrix):

Don't let your teacher hear you say "plug in -infinity", that will annoy them :) lol But yes, it looks like we're approaching this form,\[\Large\rm \frac{6}{8\cdot \infty}\]Right? So what is the expression overall approaching?

OpenStudy (el_arrow):

oh sorry so it would be 6/infinty and that equals 0 right?

zepdrix (zepdrix):

yes, the limit appears to be approaching zero overall. Good job \c:/

OpenStudy (el_arrow):

yes thank you